OpenStudy (hzro):

Use operations of decimal, fraction, and percent numbers to model the following scenarios. In your final answer, include only an equation or proportion modeling the scenario. Do not solve the problems. 5. A family dines in a popular franchise restaurant that has a policy of automatically including a 20% tip on all food and beverage sales before coupon discounts are applied. What is the family’s total bill amount if their food and beverages totaled $94.20 and they used a 15% off coupon? (HINT: Create more than one model.)

OpenStudy (sanoofy):

@MathSucks321 @mathstudent55 @amorfide

OpenStudy (cp123):

0.20x+0.15=94.20??

OpenStudy (cp123):

Thats just a guess

OpenStudy (revlogic):

intial price with tip = IP= c*1.20 cost with coupon = IP - IP*0.15 cost = 94.20 IP = c*1.20 IP = 94.20*1.20 ==> 113.04 Now, compute after using coupon: (113.04) - (113.04)(0.15) ==> 113.04 -16.956 ==> $96.08

OpenStudy (mathstudent55):

Here is a way of quickly adding or subtracting a percentage to an amount. You can use this to add a tip or subtract a discount. An item costs x. If you need to add a tip or sales tax, then think of the price of the item as 100% of its price, or 100% of x, or 1.00x. To add an 18% tip to the original price, x, the price will become 118% of the price, so now it is 1.18x. To add a 6% tax to the original price, x, the price becomes 106% of the price, or 1.06x. You can do the same with a discount or a coupon. An item costs x. x is 100% of the price of the item. The item is on sale at 20% off. That means, the price will be 100% - 20% = 80% of the original price. The discounted price, then, is 0.8x. Now let's look at your problem. price of food and drinks: x price of food and drinks plus 20% tip: 1.2x price of food and drinks plus 20% tip minus 15% coupon: 0.85(1.2x) = 1.02x Since the food and drinks cost $94.20 before the tip and the coupon, then after the tip and coupon, it will cost 1.02 * $94.20 = $96.084, which rounds off to $96.08